problem: 2x + 3y =6
how would you solve that and then graph it on a graph???

Answer by Michelle L(2)

ok hun first of you have to put it in slope formation first. that is y=mx+b k
dont worry ill walk you trough this
2x+3y=6 get y by its self
3y=-2x+b
y=-2/3x+b now you take your slope formula and plug it in with any number on a graph and it will have a negtive slope. and reamber if you are still having trouble ask a teaher that is what they are her for k